# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v4

@@ Abstract
This paper establishes a fundamental impossibility theorem: no LLM capable performing non-trivial knowledge aggregation can simultaneously achieve truthful (internally consistent) knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality. This impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses. The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity). In particular, we show how in the strictly concave settings the score of an aggregate of diverse beliefs strictly exceeds the sum of individual scores. That gap may quantify the creation of unattributable certainty or overconfidence—the mathematical origin of both hallucination and creativity, or imagination.
To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting. We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, idealized reasoning strictly preserves semantic content.
By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in the necessary violation of information conservation—this paper offers a principled foundation for managing these behaviors in advanced AI systems. Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
@@ 1 Introduction
Large language models have been (in)famous for producing convincing content that is factually incorrect, inconsistent, or well fabricated and compelling—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
@@
The idea of hallucination has clearly much to do with creativity. Therefore, we should point out at the very outset that whether it is desired or not depends on circumstances. In fact, a much better name for that phenomenon might be imagination. As an essential component of intelligence, it is responsible for seeing into the future and learning from the past. Indeed, it is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened. It is one of the driving forces of art and science, and exploration of the unknown. But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will the compelling visions of reality that does not exist.
@@
This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
@@
The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraint. We begin by modeling LLM inference as an auction of ideas where neural components—attention heads, circuits, and activation patterns—compete to influence response generation by bidding with their encoded knowledge. This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory, establishing impossibility when knowledge components are discrete and independently distributed.
@@
We then extend beyond this idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules—the standard training paradigm for modern LLMs. Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful probability reporting inevitably creates excess confidence.
@@
Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable "Jensen gap" that quantifies the violation of information conservation. This progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our theoretical choices but a fundamental property of information aggregation itself.
@@ 1.1 An Illustrative Example
Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete answer but fabricates specific details, violating truthfulness.
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response (training objectives and human preferences favor complete narratives).
@@ 1.2 The Main Result
The example above illustrates the key result of this paper that we prove formally in complementary and general settings. First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching. Second, we go beyond that idealized model and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@
For any query space $\mathcal{Q}$ containing non-trivial queries (requiring integration of available distributed knowledge), no LLM can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
@@
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge. And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
@@
In other words and on the other hand, creative tasks necessarily involve hallucination. When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation). Creativity requires controlled violation of information conservation.
Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights.
@@
Reframing the Problem: The theorem reveals that hallucination is an inherent feature of any capable inference systems. Since it is inevitable, it should be managed. Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness. Understanding these trade-offs enables principled design.
Architectural Implications: The impossibility suggests hybrid architectures where specialized modules manage different aspects of the fundamental tension. A truthfulness module might verify factual claims against known knowledge, while a creative module explores beyond the bounds of strict conservation. A meta-reasoning layer could then dynamically balance these competing outputs based on context. Training objectives should explicitly encode these trade-offs rather than naively minimizing loss across incompatible goals.
Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior. The semantic information measure quantifies knowledge accessibility within computational bounds. The emergence operator captures how reasoning transforms latent knowledge into accessible form. The Jensen gap provides a measurable signature of hallucination in actual systems.
This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself. Previous work identified important statistical correlations or architectural features associated with hallucination. We show why these patterns must exist and how to navigate possible trade-offs.
@@ 2 Related Work
Research on LLM hallucinations spans mechanistic interpretability, probabilistic frameworks, and empirical mitigation techniques. Below we present selected works in that domains. The auction-theoretic impossibility framework in this paper offers a fundamentally novel perspective distinct from these existing approaches.
@@ 2.1 Mechanistic Interpretability
Mechanistic interpretability research examines internal LLM processes to identify hallucination origins. Yu et al. yu2024mechanisticunderstandingmitigationlanguage identify two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing "agents" within LLM architectures. Meng et al. meng2023locatingeditingfactualassociations demonstrated methods for locating and editing factual associations, revealing how knowledge representations can be manipulated. Geva et al. geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes.
These findings provide concrete instantiations of our abstract "agents"—attention heads and MLP components with specialized knowledge domains competing to influence response generation. See also the fascinating and related work on the Biology of a Large Language Model lindsey2025biology .
@@ 2.2 Probabilistic Frameworks
Probabilistic approaches model hallucination through uncertainty and belief quantification. Shelmanov et al. shelmanov2025headpredictheadquestion introduced pre-trained uncertainty quantification heads for detecting hallucinations without task-specific data. Farquhar et al. farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences.
Recently, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data. While their work establishes statistical lower bounds for specific fact types under regularity assumptions, our impossibility theorem operates at a more fundamental level. We prove that any LLM facing non-trivial queries requiring knowledge integration must violate at least one of four essential properties, regardless of fact type or statistical distribution. Our framework explains why and when hallucination is inevitable, while analysis in kalai2024calibrated predicts how much hallucination to expect for facts with specific statistical properties.
@@ 2.3 Empirical Mitigation Techniques
Various empirical approaches attempt to reduce hallucination frequency with mixed success. Creswell and Shanahan creswell2022faithfulreasoningusinglarge explored faithful reasoning approaches for improving factual accuracy. Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails.
Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination—an observation our theorem explains from first principles.
Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
2.4 Theoretical Perspectives
Most significantly, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from fundamental mathematical structure, concluding that "LLMs will always hallucinate." While their conclusion aligns with that of this paper, the authors do not identify the specific trade-offs involved or provide the precise characterization our theorem offers.
3 Novel Contribution
@@
Unifying hallucination and creativity: We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
Formalizing consciousness signatures: We speculate, referring to Dennett’s ideas, that the excess confidence of response may constitute a mathematical signature of a primitive form of artificially conscious processing. That suggests consciousness might be related to sustained violations of semantic conservation by a cognitive hunger.
@@
4 Alternative Views
4.1 The Engineering Optimism View
Response: While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome. The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
4.2 The Sufficiency Threshold View
Another perspective suggests that perfect elimination of hallucination is unnecessary—sufficiently low hallucination rates could enable practical deployment without fundamental concerns. This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
Response: This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result. However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints. Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
4.3 The External Knowledge View
Response: External knowledge integration does change the game structure by altering available information during inference. However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions. While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies. The impossibility result remains independent of the ground truth available.

### v5

@@ Abstract
This paper establishes a fundamental impossibility theorem: no LLM capable of performing non-trivial knowledge aggregation can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality. The impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
We establish this result by describing the inference process as an auction of ideas, where distributed components compete exploiting their partial knowledge to shape responses. The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity). In particular, we show how to quantify the creation of overconfident or intuitive responses–the signature of both hallucination and creativity, or imagination.
To support this analysis, we introduce the complementary concepts of the semantic information measure and the emergence operator to model bounded reasoning in a general setting. We prove that while bounded reasoning generates accessible information, providing valuable insights and inspirations, the idealized unconstrained reasoning strictly preserves semantic content.
By demonstrating that hallucination and imagination are mathematically identical phenomena–grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality–we offer a principled foundation for managing these behaviors in advanced AI systems. Finally, we present some speculative ideas to inspire evaluation and refinements of the proposed theory.
@@ 1 Introduction
Large language models have been famous for producing content that is factually incorrect, inconsistent, or well fabricated and compelling–a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, like mixture-of-experts (MoE), retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
@@
The idea of hallucination has clearly much to do with creativity. Therefore, whether it is desired or not depends on circumstances. We could argue that a much better name for that phenomenon we are dealing with may be imagination. As an essential component of intelligence, imagination is responsible for seeing into the future and learning from the past. It is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened. It is one of the driving forces of art and science, and exploration of the unknown. But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions of reality that does not exist.
@@
This paper shows that hallucination is a manifestation of fundamental limitations of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality. Hallucinations will manifest one way or the other.
@@
In this paper we show the impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints. We model inference as an auction of ideas where components encoding knowledge–attention heads, circuits, and activation patterns in LLMs–compete to influence response generation by bidding with their information contributions.
@@
This game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed. We then extend beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly proper and concave scoring rules–the standard training paradigm for modern LLMs. Here, the impossibility manifests through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence. Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable (Jensen) gap that quantifies the violation of the semantic information conservation principle.
@@
The progression from abstract game theory through probabilistic frameworks to concrete neural architectures demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
@@ 1.1 An Illustrative Example
Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete and immediately satisfactory answer but fabricates specific details, violating truthfulness.
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear irritating and unsatisfying compared to the decisive but hallucinated response (training objectives and human preferences favor complete narratives).
@@ 1.2 The Main Result
The example above illustrates the core insight of this paper that we prove formally in complementary and general settings. Any response to a non-trivial query sacrifices some aspects of an ideal hallucination-free response.
First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground-truth matching. Second, we go beyond that idealized model and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
@@
No LLM can simultaneously achieve truthfulness, information conservation, knowledge revelation, and knowledge-constrained optimality when responding to non-trivial queries requiring knowledge integration.
@@
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge. And so on and so forth, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
@@
In other words and on the other hand, creative tasks necessarily involve hallucination, or better yet, imagination. When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality) or generate plausible imaginary extensions (violating conservation). Creativity requires violation of information conservation. Should we control such creative insights boldly going where no one has gone before?
@@
Despite its name, the impossibility theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights. The theorem suggests that hallucination is an inherent feature of any capable inference systems. Since it is inevitable, it should be managed. Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness. Understanding these trade-offs enables principled design.
1.3 Paper Organization
The remainder of this paper is organized as follows. Section 2 reviews the existing literature on LLM hallucinations, situating our contribution within theoretical, mechanistic, and empirical contexts. Section 3 summarizes our main contributions.
Section 4 introduces the rigorous mathematical foundations of our analysis. We define the semantic information measure ($\mu_{C}$) and the emergence operator ($\mathcal{E}_{C}$) to model bounded reasoning, establish the principle of information conservation under unlimited reasoning, and formalize the auction-theoretic model of inference. We also define the four essential properties required for hallucination-free mechanisms.
Section 5 presents the central Impossibility Theorem, offering the first two proofs. Proof I (Theorem 6) utilizes mechanism design theory in an idealized setting of independent private knowledge. Proof II (Theorem 8) generalizes the result to probabilistic settings using the theory of proper scoring rules.
Section 6 connects these theoretical results to concrete LLM implementations. We analyze the transformer architecture and present Proof III (Theorem 9), demonstrating how the Log-Sum-Exp structure inherently violates information conservation and quantifying the resulting Jensen gap ($\Gamma$).
Finally, Section 7 discusses the implications of the theorem, exploring strategies for managing trade-offs, unifying hallucination and imagination, and speculating on the connections to fundamental limits of reason and models of consciousness. Section 8 concludes the paper.
@@ 2 Related Work
Studies on LLM hallucinations span many areas of the fundamental research in AI and beyond, including neurology and philosophy. For an overview, see e.g. Huang huang25hallsurvey , Ji et al. Ji_2023 , Sun et al. Sun2024AIHT , Waters et al. Waters2016WhatIT or Boksa boksa2009neurobiology .
The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we offer here. To our best knowledge, the impossibility theorem presented in this paper offers a novel first-principles explanation for the phenomenon, distinct from existing approaches. This section situates our contribution by reviewing three key areas: theoretical arguments for the inevitability of hallucination, mechanistic studies into their origins, probabilistic frameworks for their detection, and empirical mitigation techniques.
2.1 Theoretical Perspectives on Inevitability
In their work, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from their fundamental mathematical and logical structure. Their conclusion is that LLMs cannot detect and eliminate hallucinations by themselves and therefore must hallucinate. The analysis draws on Goedel’s incompleteness theorem and Turing’s undecidability theorem. First, incompleteness argument is called to show that an LLM cannot verify its responses because of the training set incompleteness. Second, the hallucination (self)detection problem is reduced to the halting problem, which is undecidable. While the finale conclusion aligns with that of our paper, the arguments presented are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations. In other words, it is about practical impossibility of real-time, exhaustive fact-checking.
Similar arguments are presented in other works. Xu et al. Xu2024HallucinationII define hallucination as inconsistency between a computable LLM and a computable ground truth function. Then the paper shows that LLMs cannot learn all the computable functions and will therefore inevitably hallucinate. Suzuki et al. suzuki2025hallucinations show that hallucinations can be made statistically negligible, provided that the quality and quantity of the training data are sufficient. Then, Wu et al. show in Wu2024NoFL that non-hallucinating learning is statistically impossible when relying solely on the training dataset.
Our impossibility theorem goes beyond those approaches operating at a more fundamental level of the inference process itself.
@@ 2.2 Mechanistic Interpretability
Mechanistic interpretability examines internal LLM processes to identify emergence of ideas, interpretable concepts and dynamics of their interactions, including hallucination origins, see Sharkey et al. Sharkey2025OpenPI for an overview.
Yu et al. yu2024mechanisticunderstandingmitigationlanguage identified signatures in model internals that predict a likelihood of hallucinating. The paper identifies two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing agents within LLM architectures. Meng et al. meng2023locatingeditingfactualassociations demonstrated methods for locating and editing factual associations, revealing how knowledge representations can be manipulated. Geva et al. geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes. More recently, Gottesman and Geva Gottesman2024EstimatingKI have proposed a method to estimate how knowledgeable a model is about a certain entity and identify tokens indicative of clusters and gaps in the model’s knowledge.
These findings provide concrete instantiations of our abstract agents–attention heads and MLP components with specialized knowledge domains competing to influence response generation. We must refer here to the fascinating and related work of Lindsey et al. lindsey2025biology on the biology of a Large Language Model. It is an investigation of the internal chains of steps that a model uses to transform a specific input prompt into an output response. In particular, the paper discovers how LLM plans its outputs ahead of time when writing lines of poetry, identifying in advance potential rhyming words that could appear at the end. We can view that as an empirical illustration of our auction of ideas.
@@ 2.3 Probabilistic Frameworks
Probabilistic approaches model and asses hallucination through uncertainty and belief quantification. Shelmanov et al. shelmanov2025headpredictheadquestion introduced pre-trained uncertainty quantification heads for detecting hallucinations without task-specific data.
Farquhar et al. farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences. That concept of semantic entropy is well aligned and complementary to our idea of semantic information measure and semantic independence, pointing out the importance of understanding the meaning of competing ideas in the studies of imagination.
Most notably, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data. While that work establishes statistical lower bounds for specific types of hallucination, our impossibility theorem operates at a more fundamental and general level. We prove that any LLM facing non-trivial queries requiring knowledge integration must violate at least one of four essential properties, regardless of fact type or statistical distribution. Our framework explains why and when hallucination is inevitable, while analysis in kalai2024calibrated predicts how much hallucination to expect for facts with specific statistical properties.
@@ 2.4 Empirical Mitigation Techniques
Various empirical approaches attempt to reduce hallucination frequency with mixed success. For a recent survey see e.g. Tonmoy et al Tonmoy2024ACS . Here we refer only to selected works.
Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails. Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination–an observation our theorem explains from first principles. Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
3 Main Contributions
@@
Unifying hallucination and creativity (imagination): We prove these are mathematically identical phenomena that arise from the convex aggregation process that creates confidence exceeding constituent evidence.
@@
Theoretical Foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior. The semantic information measure quantifies knowledge accessibility within computational bounds. The emergence operator captures how reasoning transforms latent knowledge into accessible form. The Jensen gap provides a measurable signature of hallucination in actual systems. This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself. Previous work identified important statistical correlations or architectural features associated with hallucination. We show why these patterns must exist and how to navigate possible trade-offs.